# Sparsity May Be All You Need: Sparse Random Parameter Adaptation

### v1

@@ 7.2 IMDB
gemma-2b gemma-2b-it mistral-7b mistral-7b-it Adaptation density (%) loss acc. loss acc. loss acc. loss acc. Full FT 100% 0.092 96.9% 0.107 96.3% 0.080 97.4% 0.080 97.3% SpaRTA 5% 0.096 96.7% 0.105 96.3% 0.084 97.2% 0.081 97.5% SpaRTA 0.5% 0.096 96.7% 0.103 96.4% 0.087 96.9% 0.087 97.0% SpaRTA 0.05% 0.106 96.3% 0.114 96.2% 0.080 97.3% 0.076 97.4% LoRA $\approx$0.050.101 96.5% 0.113 96.2% 0.086 97.1% 0.081 97.1% Head $\approx$2e-40.165 93.6% 0.194 92.7% 0.139 94.9% 0.117 95.8%
@@
Table 3: Test loss and accuracy of models adapted to the IMDB dataset with different fine-tuning methods. For training details see Table 12.
Table 3 shows the results for the IMDB classification task where each model is asked to rate a review as positive or negative. Each model is fine-tuned using adaptation techniques including: (i) Full FT where all model parameters are updated (100% density); (ii) SpaRTA for different orders of magnitude of density levels 5%, 0.5%, $\approx$0.05– the last one allowing for SpaRTA to have close to the same number of trainable parameters as for LoRA; (iii) LoRA with rank $r=8$ equivalent to $\approx 0.05\%$ of trainable parameters compared to the model full parameter size; (iv) Head adaptation where only the classification head is updated ($\approx$2e-4density).
@@
Model: gemma-2-2b-it QNLI RTE SST2 QQP MNLI MRPC COLA loss acc. loss acc. loss acc. loss acc. loss acc. loss acc. loss mcc Full FT (100%) 0.46 81.2 0.33 87.4 0.18 93.3 0.36 85.5 0.26 88.9 0.12 96.1 0.36 63.6 SpaRTA (5%) 0.45 79.8 0.33 87.6 0.18 93.0 0.40 81.3 0.24 89.9 0.21 95.6 0.41 57.2 SpaRTA (0.5%) 0.55 72.9 0.37 86.1 0.25 89.8 0.34 85.4 0.24 89.9 0.13 95.6 0.39 56.9 SpaRTA (0.037%) 0.45 79.8 0.38 85.5 0.18 93.0 0.40 82.3 0.27 88.6 0.12 95.8 0.45 56.5 LoRA (0.037%) 0.46 78.3 0.33 87.5 0.24 90.4 0.36 84.8 0.26 89.0 0.13 95.5 0.39 59.9 Head (1.63e-4%) 2.60 47.3 0.73 67.4 0.96 37.1 0.55 74.2 0.66 64.9 0.55 84.2 0.62 18.0
@@
Table 4: Test loss and accuracy of model google/gemma-2-2b-it adapted to glue datasets with different fine-tuning methods. For training details see Table 13.
@@
In Table 3, the results for adaptation techniques (rows) are sorted in order of descending density as to show the impact of decreasing the number of trainable parameters on the overall performance. For gemma-2b and gemma-2b-it, Full FT adaptation gives strong test loss and accuracy numbers, often the best available – this is expected since all model parameters are fine-tuned. SpaRTA results at 5% density are close to Full FT, even matching them for gemma-2b-it. As the density decreases by orders of magnitude, the results slowly degrade. For $\approx$0.05%, gemma models’ performances are slightly behind or matching performance from LoRA. Head adaptation provides the worse results as one would expect since it has so few parameters to work with.
@@
For mistral models, the trend is similar with Full FT adaptation showing strong loss and accuracy numbers. SpaRTA performs well, matching and even improving upon Full FT with a density of 0.05%. SpaRTA even improves over LoRA for both mistral-7b and mistral-7b-it at this low density. Once again, Head adaptation gives the worst results. Overall these results are encouraging and show that SpaRTA is competitive and provides similar performance to LoRA for low densities.
7.3 GLUE
@@
Model: Mistral-7B-Instruct-v0.3 QNLI RTE SST2 QQP MNLI MRPC COLA loss acc. loss acc. loss acc. loss acc. loss acc. loss acc. loss mcc Full FT (100%) 0.32 91.7 0.26 91.3 0.14 95.7 0.33 88.6 0.23 90.3 0.11 96.7 0.32 70.5 SpaRTA (5%) 0.30 89.9 0.23 91.4 0.12 95.8 0.28 88.9 0.23 90.5 0.10 96.7 0.30 72.1 SpaRTA (0.5%) 0.29 89.5 0.23 91.6 0.13 95.2 0.29 88.6 0.25 89.2 0.10 96.6 0.32 70.3 SpaRTA (0.048%) 0.27 89.9 0.24 91.1 0.12 95.9 0.31 87.6 0.25 89.3 0.10 96.6 0.32 69.9 LoRA (0.048%) 0.32 88.8 0.26 90.4 0.12 95.6 0.32 87.2 0.27 88.5 0.11 96.9 0.30 71.3 Head (1.15e-4%) 0.31 86.3 0.35 87.0 0.23 90.4 0.35 84.5 0.32 85.5 0.15 93.5 0.36 62.1
@@
Table 5: Test loss and accuracy of the mistralai/Mistral-7B-Instruct-v0.3 model adapted to glue datasets with different fine-tuning methods. For training details see Table 13.
Table 4 compares the results of gemma-2b-it on 7 of the GLUE tasks for the same adaptation methods as in Table 3. An equivalent set of results for the gemma-2b base model is given in Table 10, see Appendix D. Here again, adaptation results (rows) are ordered in descending order of density with SpaRTA evaluated for different orders of magnitude of density levels 5%, 0.5%, 0.037% – the last one allowing SpaRTA to have approximately the same number of trainable parameters as LoRA. LoRA has rank $r=8$ equivalent to 0.037% of trainable parameters; Head adaptation uses a very small 1.63e-4% density. Overall, the same trend as for IMDB is observed. Full FT results are very often the best, sometimes bested by SpaRTA 5% (RTE, MNLI). For a low density, where LoRA and SpaRTA use the same number of trainable parameters, SpaRTA can best LoRA (QNLI, SST2, RTE) but is overall competitive with LoRA results.
@@
A similar set of results on the same GLUE tasks is provided for mistral-7b-it in Table 5. Again, results for the mistral-7b PT base model can be found in Table 11, see Appendix D. Full FT results can often be matched or bested by SpaRTA 5% (RTE, SST2, QQP, MNLI, MRPC, COLA). SpaRTA often gives better results than LoRA (QNLI, RTE, SST2, QQP, MNLI) for a low density (0.0048%) where both techniques have comparable numbers of parameters. Once again, SpaRTA is a competitive adaptation technique, often providing similar results to LoRA with low density.
@@ 7.4 Remarks
Results on GLUE and IMDB establish that SpaRTA is a viable adaptation technique that can be competitive with LoRA, especially for larger LMs. These results demonstrate that a simple sparsifying scheme can offer a valid adaptation technique. This opens up the possibility for more investigation in sparse adaptation as the low rank approximation of a model’s parameter changes is not the only mechanism to provide a performant adaptation method. This indicates that what matters is not necessarily the adapter structure used in PEFT techniques, but rather the number of independent parameters, that is relevant to the adaptation task at hand.
8 Exploring Model Merging
Yu et al. (2024) demonstrate that one can perform Full FT, compute $\Delta=\theta_{\text{FT}}-\theta_{\text{PT}}$, sparsify $\Delta$ to get $\tilde{\Delta}$, and obtain good performance using $\tilde{\theta}_{\text{FT}}=\theta_{\text{PT}}+\tilde{\Delta}$ as model parameters. Furthermore, they show that merging models via their sparsified $\tilde{\Delta}$-s from different tasks can also maintain the performance of the individual models. This motivates exploring model merging via SpaRTA, where no Full FT is required and sparse adapters are instead fine-tuned.
In our setup, because the heads of our SpaRTA-tuned models are fully trained, merging them into a single head would create parameter interference. We thus merge everything except the heads and add each unmerged head on top of the merged model creating a multi-head model. We need the merged model to produce multiple outputs (i.e., one for each task) per input and this is exactly what a multi-head merged model does. In the forward pass, a sequence of input tokens goes through the merged model, producing a final state representation of the tokens. This representation is shared by all tasks and is passed to each head to produce one prediction per task as output. This way all tasks are processed concurrently on the same input.
We experiment merging two models: one individually fine-tuned with SpaRTA for SST2 (classifying text sentiment) and another for COLA (classifying grammatical correctness), using gemma-2b as PT model and 99% sparsity in both cases. We merge these two models by adding the respective $\Delta_{\phi}$ to their PT model’s original parameters $\theta_{\text{PT}}$, and obtaining a two-headed model that can simultaneously classify any input sentence on both criteria in a single forward pass. Table 6 compares performance of the merged model with that of the unmerged models. This result encourages future exploration of this practical usecase for SpaRTA.
SST2 COLA
Model loss acc loss acc
SST2 0.113 96.3% - -
COLA - - 0.487 82.7%
Merged 0.118 96.7% 0.428 84.4%
Table 6: Test loss an accuracy of SST2 and COLA models fine-tuned on gemma-2b using SpaRTA with 99% sparsity. The merged model outperforms both individual models in this scenario.
@@ 9 Conclusion
As PT language models have grown in size, PEFT has become crucial for enabling fine-tuning large PT language models on limited hardware and financial resources. We have introduced SpaRTA, an approach that sharply decreases the set of trainable parameters, reducing the GPU memory used by the optimizer and speeding up training. We have demonstrated on a variety of task adaptation scenarios that our fine-tuning approach is parameter-efficient and competitive with LoRA, the current PEFT standard. Experiments with 2B and 7B parameter pre-trained models demonstrate good performance, and as per Yu et al. (2024), we expect larger models to allow for higher levels of sparsity in training; meaning that efficiency of SpaRTA should get better with larger model sizes as also suggested in Table 1.
Regarding future directions, while SpaRTA has been applied to supervised learning, it is also amenable to reinforcement learning often used for model alignment (Ouyang et al., 2022a). Furthermore, SpaRTA allows one to select which modules to sparsify and at what layers. Like LoRA, this decision affects both the number of trainable parameters and the performance of the fine-tuned model, and we plan to investigate its impact in SpaRTA. Lastly, we plan to explore model merging as discussed in Section 8; SpaRTA opens the potential for model merging without Full FT but with possibly little interference.
@@ Appendix A Ranks of differences in fine-tuned weight matrices
We compute all delta weight matrices for two well-known instruction-following fine-tuned models: gemma-2b-it and mistral-7b-it, see Section 6.2. As we can see in Tables 7 and 8, the feedforward (MLP) delta $\Delta$-matrices associated with each layer of the transformer network are all full rank in both models. This is also the case for all self attention key (k) and value (v) projection matrices. The self attention query (q) and output (o) $\Delta$-matrices all show relatively small rank deficiencies: between $9$ and $2$ for the query projection matrices and between $46$ and $3$ for the output projection matrices, out of a potential maximum rank of $2,048$ for the gemma-2b-it model; and between $1,788$ and $28$ for the query projection matrices and between $246$ and $10$ for the output projection matrices, out of a potential maximum rank of $4,096$ for mistral-7b-it model. The token embedding $\Delta$-matrix for mistral-7b-it is full rank. However, it has a rank of $97$ for gemma-2b-it, being the only matrix for that model whose delta change shows a significant rank deficiency (of $1,951$). Of the two models, mistral-7b-it is the only one that does not tie its tokens embedding weights to its head; so that mistral-7b-it model’s head (untied LM head), which is trained independently from the tokens embedding, has a rank deficiency of $1$ for its delta change. Basically, we observe that the fine-tuning changes in all weight matrices of these models are full rank or very close to it, with the only exception being the changes in the tokens embedding for gemma-2b-it which are significantly low rank.
We analyzed here the changes in fine-tuned weight matrices for two popular open-weights language models. As discussed in Section 2.1, we compared instruction-following models with their original, pre-trained versions, by computing the differences (delta) in their weight matrices. Tables 7 and 8 provide a detailed breakdown of the ranks for all these delta matrices. The observed ranks suggest that constraining the (delta) changes in weight matrices to be low-rank is not essential for fine-tuning models efficiently.
Weight matrix dims Layer(s) Rank Rank deficiency
tokens embed [ 256000, 2048 ] - 97 1951
self attn q proj [ 2048, 2048 ] 0 2039 9
15 2041 7
14 2042 6
6 2044 4
1,3-5,7,8,10-13,16,17 2045 3
2,9 2046 2
@@
self attn o proj [ 2048, 2048 ] 0 2002 46
16 2040 8
11,12,14 2042 6
6,8,10,13,15 2043 5
3-5,7,9,17 2044 4
1,2 2045 3
@@
Weight matrix dims Layer(s) Rank Rank deficiency
@@
self attn q proj [ 4096, 4096 ] 0 2308 1788
1 3103 993
3 4023 73
31 4046 50
30 4050 46
2,4 4051 45
6 4052 44
29 4056 40
28 4057 39
11,27 4058 38
8,9 4059 37
10,19,25,26 4061 35
12,18,23 4062 34
5,7,14,17 4063 33
13,20,22 4064 32
15 4065 31
21,24 4066 30
16 4068 28
@@
self attn o proj [ 4096, 4096 ] 0 3850 246
31 4059 37
1,30 4060 36
14 4068 28
18 4069 27
23 4070 26
2,10,26 4071 25
8,27 4072 24
3,9 4074 22
4,11,29 4075 21
6,13,22,28 4076 20
7,12 4077 19
15,16,19 4079 17
17 4080 16
5 4081 15
25 4083 13
21,24 4085 11
20 4086 10
@@
untied lm head [ 4096, 32768 ] - 4095 1

### v2

@@ 7.1 IMDB
Sparsification targets Loss Accuracy
Wq, Wv (like LoRA) 0.117 96.3%
Wv, Wo 0.109 96.7%
Wq, Wk, Wv 0.123 95.9%
Wq, Wk, Wo 0.112 96.0%
Wq, Wq, Wv, Wo 0.110 96.5%
MLP 0.150 94.7%
Wq, MLP 0.150 94.5%
Wk, MLP 0.150 95.1%
Wv, MLP 0.145 95.2%
Wo, MLP 0.140 95.6%
Wq, Wk, MLP 0.150 94.6%
Wv, Wo, MLP 0.136 95.8%
W, MLP, norm 0.140 95.4%
Table 3: Full table of loss and accuracy on the SST-2 task after applying our SpaRTA approach to different types of parameters in the Gemma 2B model, given that the same number of trainable parameters are selected. Results are averaged across 10 random seeds
@@
gemma-2b gemma-2b-it mistral-7b mistral-7b-it method density (%) loss acc. loss acc. loss acc. loss acc. PT (zero-shot) 0.425 92.0% 0.435 86.1% Full FT 100% 0.092 96.9% 0.107 96.3% 0.080 97.4% 0.080 97.3% SpaRTA 5% 0.096 96.7% 0.105 96.3% 0.084 97.2% 0.081 97.5% SpaRTA 0.5% 0.096 96.7% 0.103 96.4% 0.087 96.9% 0.087 97.0% SpaRTA 0.05% 0.106 96.3% 0.114 96.2% 0.080 97.3% 0.076 97.4% LoRA $\approx$0.050.101 96.5% 0.113 96.2% 0.086 97.1% 0.081 97.1%
@@
Table 4: Test loss and accuracy of models adapted to the IMDB dataset with different fine-tuning methods. We also report zero-shot performance of instruction following PT models. For training details see Table 16.
@@
Model: gemma-2b-it QNLI RTE SST2 QQP MNLI MRPC COLA BoolQ method targets loss acc. loss acc. loss acc. loss acc. loss acc. loss acc. loss mcc loss acc. avg. PT (zero-shot) 1.34 40.5 1.77 43.7 0.75 60.7 1.41 48.3 2.91 30.9 0.91 64.6 0.85 0.7 0.88 68.5 44.7 SpaRTA (5%) ALL 0.17 93.7 0.54 81.0 0.14 95.0 0.26 89.0 0.33 87.2 0.35 85.3 0.39 56.2 0.36 84.4 84.0 SpaRTA (0.5%) 0.16 94.0 0.43 80.7 0.15 94.7 0.24 89.7 0.32 87.8 0.34 86.5 0.45 54.9 0.37 84.7 84.1 SpaRTA (0.037%) 0.18 93.1 0.48 76.9 0.16 94.5 0.26 88.8 0.35 86.4 0.36 84.5 0.45 54.8 0.40 83.6 82.8 LoRA (0.037%) Q,V 0.18 93.4 0.44 78.7 0.14 95.1 0.26 89.1 0.33 87.8 0.33 85.4 0.41 55.0 0.36 84.0 83.6 DoRA (0.037%) 0.18 93.4 0.45 78.0 0.15 95.6 0.26 89.0 0.33 87.8 0.33 85.5 0.41 57.1 0.36 84.4 83.8 SpaRTA (0.037%) 0.19 92.8 0.46 77.6 0.16 94.8 0.26 88.6 0.36 86.3 0.36 84.2 0.45 54.5 0.39 83.4 82.8 SpaRTA (0.037%) O,V 0.17 93.5 0.44 80.7 0.15 94.9 0.26 88.9 0.34 87.2 0.35 85.7 0.42 55.9 0.37 85.0 84.0
@@
Table 5: Test loss and accuracy of model gemma-2b-it adapted to GLUE and BoolQ datasets with different fine-tuning methods. For training details see Appendix E.2.
Table 4 shows the results for IMDB, where each model is asked to classify a review as positive or negative. Each model is fine-tuned using the following adaptation methods: (i) Full parameter fine-tuning (Full FT) where all model parameters are optimized; (ii) SpaRTA for different density levels 5%, 0.5%, 0.05%, with the last allowing SpaRTA to have approximately the same number of trainable parameters as LoRA; (iii) LoRA with rank $r=8$ equivalent to about $0.05\%$ of trainable parameters compared to the model full parameter size.
@@
In Table 4, results for adaptation methods (rows) are sorted in order of descending density as to show the impact of decreasing the number of trainable parameters on the overall performance. For gemma-2b and gemma-2b-it, Full FT adaptation gives (practically always) the best test loss and accuracy numbers. This is expected since all model parameters are fine-tuned. SpaRTA results at 5% density are close to Full FT, even matching them for gemma-2b-it. As the density decreases by orders of magnitude, the results slowly degrade. For 0.05% density, gemma models’ performances match or slightly lag performance from LoRA.
@@
For mistral models, the trend is similar with Full FT showing best (or close to best) loss and accuracy numbers. SpaRTA performs well, matching and even improving upon Full FT with a density of 0.05%. SpaRTA even improves over LoRA for both mistral-7b and mistral-7b-it at this low density. Overall these results are encouraging; SpaRTA is competitive, provides similar performance to LoRA for low densities, and does not degrade the performance of full parameter fine-tuning.
7.2 Which parameters should SpaRTA target?
So far, SpaRTA has chosen trainable parameters from any layer of the transformer with equal probability. In contrast, LoRA concentrates adaption on the key and value self-attention weight matrices.
Now, we investigate whether it is better to concentrate the selection of sparse, trainable parameters on a few parameters or equally across all. Specifically, given a budget for the number of parameters that can be trained, which parameters should we target for sparsification (while freezing the remaining) to achieve best task performance when adapting the sparsified parameters within the targeted types?
We conduct an ablation study to answer this question. The Gemma 2B model is adapted to the SST-2 task using our SpaRTA approach targeting different combinations of parameters. We set a budget of 1.25M trainable parameters within the Gemma 2B model so that sparsity $s=99.95\%$.
Results of these experiments are shown in Table 3. Concentrating the selection of trainable parameters through sparsification in the self-attention value (Wv) or/and output (Wo) weight matrices yields the best performance under the given budget. In contrast, distributing our sparsification across different combinations of weight types may lead to significantly lower performance.
Note that concentrating our sparsification over targeted parameters, although proven beneficial in terms of single-task adaption performance, can decrease performance when merging since it increases the chance of parameter interference.
7.3 GLUE and BoolQ
@@
Model: Mistral-7B-Instruct-v0.3 QNLI RTE SST2 QQP MNLI MRPC COLA BoolQ method targets loss acc. loss acc. loss acc. loss acc. loss acc. loss acc. loss mcc loss acc. avg. PT (zero-shot) 2.38 23.4 2.54 33.6 0.66 66.3 0.85 57.5 2.42 35.5 0.70 70.0 0.72 25.2 0.39 85.2 49.6 SpaRTA (5%) ALL 0.11 96.0 0.26 89.8 0.12 95.9 0.22 90.7 0.23 91.6 0.32 87.3 0.36 62.6 0.24 90.8 88.1 SpaRTA (0.5%) 0.11 95.8 0.30 90.5 0.12 96.0 0.23 90.4 0.24 91.2 0.32 88.0 0.35 67.9 0.24 90.5 88.8 SpaRTA (0.048%) 0.12 95.7 0.31 89.2 0.12 96.1 0.23 90.5 0.24 91.1 0.31 88.2 0.34 67.3 0.26 89.5 88.4 LoRA (0.048%) Q,V 0.12 95.6 0.24 91.2 0.12 96.0 0.23 90.2 0.24 91.0 0.29 89.3 0.32 69.9 0.24 90.5 89.2 DoRA (0.048%) 0.13 95.0 0.27 90.9 0.11 96.4 0.23 90.2 0.25 90.9 0.28 88.7 0.32 68.2 0.24 90.5 88.8 SpaRTA (0.048%) 0.12 95.6 0.31 88.3 0.13 95.6 0.23 90.1 0.25 90.9 0.30 88.6 0.35 65.2 0.26 89.7 88.0 SpaRTA (0.048%) O,V 0.11 95.9 0.31 88.4 0.14 95.8 0.23 90.4 0.24 91.2 0.30 88.4 0.34 64.9 0.26 89.7 88.1
@@
Table 6: Test loss and accuracy of the mistral-7b-it model adapted to GLUE and BoolQ datasets with different fine-tuning methods. For training details see Appendix E.2.
MMLU gemma-2b gemma-2b-it mistral-7b mistral-7b-it loss acc. loss acc. loss acc. loss acc. PT (zero-shot) - - 5.284 35.3% - - 1.839 59.3% SpaRTA 1.249 45.1% 1.250 45.1% 0.987 61.5% 0.930 63.1% LoRA 1.233 46.9% 1.271 45.1% 0.981 62.8% 0.928 63.0%
Table 7: Test loss and accuracy of models adapted to the MMLU dataset with SpaRTA and LoRA. We also report zero-shot performance of instruction following PT models. For training details see Table 17 in the Appendix.
We now focus on comparing SpaRTA and LoRA methods using the 7 NLU task in GLUE tasks as well as BoolQ. Table 5 presents the results for adapting gemma-2b-it. An equivalent set of results for the gemma-2b base model is given in Table 10, see Appendix D. Adaptation results (rows) are grouped by which type of parameters are targeted, with SpaRTA further ordered in descending order of density. SpaRTA with 0.037% density has approximately the same number of trainable parameters as LoRA, for which a rank $r=8$ was used.
When targeting all parameters, SpaRTA overall exhibits better performance with higher density, i.e, when more parameters are chosen to be trained, with a few exceptions. Targeting the Wq and Wv self-attention matrices shows very similar performance between LoRA, DoRA, and SpaRTA. Targeting the Wo and Wv self-attention matrices within SpaRTA, which was shown to be the optimal in Section 7.2, yields better performance for SpaRTA when compared to both LoRA and DoRA on four of the eight datasets (QNLI, RTE, MRPC, BoolQ).
@@
A similar set of results on GLUE and BoolQ is provided for mistral-7b-it in Table 6. Again, results for the mistral-7b base model can be found in Table 11 in Appendix D. Again, SpaRTA exhibits comparable performance to LoRA and DoRA for a low density of 0.0048%, where all methods have comparable numbers of trainable parameters. And once again, with SpaRTA targeting Wv and Wo typically outperforming the targeting of Wq and Wv.
7.4 MMLU
We have chosen MMLU because it is known for being a very challenging NLU task. Table 7 compares SpaRTA against LoRA on adapting each of our models to the MMLU multiple choice question answering task, where each model must predict the correct answer to a set of questions. Solving this task requires a high level of world knowledge and problem solving skills. In this experiment, we restrict both methods to use approximately the same number of trainable parameters. See Table 17 in the Appendix for training details. Here again, the results show that SpaRTA is competitive with LoRA, providing similar performance.
@@ 7.5 Remarks
Our results establish that SpaRTA is a viable adaptation technique that can be competitive with LoRA, especially for larger LMs. These results demonstrate that a simple sparsifying scheme can offer a valid adaptation technique. This opens the possibility for further investigating sparse adaptation as the low rank approximation of a model’s parameter changes is not the only mechanism to provide a performant adaptation method. This indicates that what matters is not necessarily the adapter structure used in PEFT techniques, but rather the number of trainable parameters, that is relevant to the adaptation task at hand.
@@ 8 Conclusion
As PT language models have grown in size, PEFT has become crucial for enabling fine-tuning large PT language models on limited hardware and financial resources. We have introduced SpaRTA, an approach that sharply decreases the set of trainable parameters, reducing the GPU memory used by the optimizer and speeding up training. We have demonstrated on a variety of task adaptation scenarios that our fine-tuning approach is parameter-efficient and competitive with LoRA, the current PEFT standard. Experiments with 2B and 7B parameter pre-trained models demonstrate good performance, and as per Yu et al. (2024), we expect larger models to allow for higher levels of sparsity in training, meaning that efficiency of SpaRTA should get better with larger model sizes as also suggested in Table 1.
Regarding future directions, while SpaRTA has been applied to supervised learning, it is also amenable to reinforcement learning often used for model alignment (Ouyang et al., 2022). We also plan to explore the merging of SpaRTA adapters due to their potential for little interference.
@@ Appendix A Ranks of differences in fine-tuned weight matrices
We compute all delta weight matrices for two well-known instruction-following fine-tuned models: gemma-2b-it and mistral-7b-it, see Section 6.2. As we can see in Tables 8 and 9, the feedforward (MLP) delta $\Delta$-matrices associated with each layer of the transformer network are all full rank [^5] We compute the rank of a matrix as the number of singular values that are greater than zero, given a specified tolerance, using torch.linalg.matrix_rank. We did some sensitivity analysis to determine such tolerance. in both models. This is also the case for all self attention key (k) and value (v) projection matrices. However, the self attention query (q) and output (o) $\Delta$-matrices all show relatively small rank deficiencies: between $9$ and $2$ for the query projection matrices and between $46$ and $3$ for the output projection matrices, out of a potential maximum rank of $2,048$ for the gemma-2b-it model; and between $1,788$ and $28$ for the query projection matrices and between $246$ and $10$ for the output projection matrices, out of a potential maximum rank of $4,096$ for the mistral-7b-it model. The token embedding $\Delta$-matrices of both the gemma-2b-it and mistral-7b-it models are full rank. Of the two models, mistral-7b-it is the only one that does not tie its token embeddings’ weights to its head; thus the mistral-7b-it’s model head (untied lm head) is updated independently during fine-tuning, and the resulting delta change of this matrix is also full rank. Basically, we observe that the fine-tuning changes in the weight matrices of these well known models are all full rank, with the only exception being the changes in the query (q) and output (o) projection matrices that show small rank deficiencies.
The observed ranks suggest that constraining the (delta) changes in weight matrices to be low-rank is not essential for fine-tuning models efficiently.
Weight matrix dims Layer(s) Rank(s) Rank deficiencies
tokens embed [ 256000, 2048 ] - 2048 0 (full rank)
self attn q proj [ 2048, 2048 ] 0 - 17 2046 - 2039 2 - 9
@@
self attn o proj [ 2048, 2048 ] 0 - 17 2045 - 2002 3 - 46
@@
Weight matrix dims Layer(s) Rank(s) Rank deficiencies
@@
self attn q proj [ 4096, 4096 ] 0 - 31 4068 - 2308 28 - 1788
@@
self attn o proj [ 4096, 4096 ] 0 - 31 4086 - 3850 10 - 246
@@
untied lm head [ 4096, 32768 ] - 4096 0 (full rank)